# Navigator: A Decentralized Scheduler for Latency-Sensitive ML Workflows

### 27 Feb 2024 v1

@@ Abstract
We consider ML query processing in distributed systems where GPU-enabled workers coordinate to execute complex queries: a computing style often seen in applications that interact with users in support of image processing and natural language processing. In such systems, coscheduling of GPU memory management and task placement represents a promising opportunity. We propose Navigator, a novel framework that unifies these functions to reduce job latency while using resources efficiently, placing tasks where data dependencies will be satisfied, collocating tasks from the same job (when this will not overload the host or its GPU), and efficiently managing GPU memory. Comparison with other state of the art schedulers shows a significant reduction in completion times while requiring the same amount or even fewer resources. In one case, just half the servers were needed for processing the same workload.
@@ 1 Introduction
The work described here is motivated by a recent trend: under pressure to respond rapidly to incoming events, the cloud is expanding to include edge clusters physically close to the end-user. Our new scheduler, Navigator, targets this setting, managing resources in ways that improve job completion delay with the same (or reduced) level of resources. Machine intelligence is already popular in interactive applications, and projected to gain importance for at least another decade, so progress could be broadly valuable.
@@
Our work represents edge ML applications as acyclic data-flow graphs (DFGs). Each ML instance runs in its own address space hence a single application could include ML pipelines implemented using different tools (we currently support PyTorch [34], TensorFlow [19], and MXNet [45]). A triggering event launches one job, and each DFG step within it runs as a task on a single server (we will relax these restrictions in future work). Navigator plays two roles: platform-level GPU cache management and job/task placement. Our algorithm has the primary goal of minimizing job latency, with secondary goals of achieving efficient hardware utilization and packing work into as few servers as feasible so that unused hardware can be put into a power-saving mode.
An important decision was to make Navigator fully decentralized and symmetric: every worker node can schedule tasks on every other worker. While many existing schedulers [36, 53, 50, 47, 64, 48] are centralized, permitting fully coordinated decision-making, centralization introduces overheads due to the need to consult the scheduler and to notify it of changing conditions, creating possible bottlenecks. Early in our effort, we realized that in edge systems task execution times are not fully predictable, hence worker loads and GPU memory contents can quickly evolve in unanticipated ways. Decentralized tracking of system state was traditionally assumed to be far more costly, but modern data replication is inexpensive and scalable [37]. Our decentralized solution has low overheads and outperforms centralized alternatives.
@@
    Experiments and simulation showing that Navigator can reduce latency by 2x-6x. Navigator never consumes more resources than existing schedulers and requires as few as half the servers to run an identical workload.
@@ 2.1 Dataflow Graphs
Navigator’s DFG representation is similar to that used in prior work [10, 16, 74, 53]. As noted earlier, these are directed acyclic graphs $G=(E,V)$, where vertex $v\in V$ represents an ML computation. Edge $e\in E$ represents precedence constraints: output from the upstream ML will become input to the downstream ML. We attach a diamond box to vertex $v$ to represent data dependencies: the ML model and other objects that $v$ requires to perform its computational task. As seen in Figure 1, identical colors denote the same ML model, while different ones denote different models. Our figures omit details such as object sizes and estimated job/task execution times, but the DFG also includes them. Our experimental section used these workflows as its scheduling target.
@@ 2.2 Deployment Assumptions
Navigator could be used in a cloud datacenter, but will often be deployed in edge clusters located near data-capture devices, an approach also adopted in prior work  [69]. These clusters would be more limited than full-fledged cloud servers but still could host multiple DFGs to serve mixes of edge requests. Using secure cloud links, edge applications would handle time-sensitive tasks while drawing on the full cloud for aspects off the low-latency critical path.
@@ 2.3 Scheduler Objectives
Our four example ML workflows all require fast responses. For example, when performing language translation and image generation, end-to-end performance of the DFG determines the wait time the human user will perceive. Navigator’s primary role is to minimize end-to-end latency, defined as the difference between the end time (when the processing of the exit task is finished) and the start time (when the job instance is generated). Minimizing latency in DFG structured workflows is NP-complete, hence Navigator will inevitably depend on heuristics [9, 11, 7]. We do not explicitly optimize for resource consumption or energy, but our algorithm turns out to be parsimonious in both respects.
@@ 3 System Architecture
As shown in Figure 2, we assume that each worker consists of a host supporting storage, host computation, and a GPU unit for heavy lifting. All workers run Navigator, which is a fully decentralized and symmetric: every worker has visibility of the system state and can assign tasks to any of its peers.
Figure 2: Worker Components in Navigator
@@
The architecture of Navigator is seen in Figure 2 and consists of the following components: Workflow Profiling collects profiled information about the vertices in a DFG such as average execution times. The Decentralized Global State Monitor collects and periodically exchanges information about all the workers: their load, cache bitmap, etc. Navigator’s Scheduler component creates a ADFG and later adjusts it. The task dispatcher executes the decisions taken by the Scheduler, sending task start requests to the correct worker as the prior task(s) complete. The GPU Memory Manager fetches models on the fly and manages a GPU cache. Task processing is managed by the Execution Engine, which includes a plug-in customized for each supported ML framework. We now describe each of the components, while the details of Navigator’s scheduling algorithm are presented in Section 4.
@@ 3.1 Repository of Workflow Profiles
Navigator’s Workflow Profiles Repository holds meta-information about DFGs. In addition to static information (expected runtime costs and input/output object sizes), each task is annotated with the actual sizes of known inputs.
@@ 3.2 Scheduler and Task Dispatcher
Preliminary scheduling decisions enable Navigator to collocate a series of tasks on the same worker. This saves time by avoiding the need to move the output of one task to the location where the next worker would run. Moreover some aspects of task assignment can only occur at the outset: if a set of tasks are followed by a join, they would have no way to make a coordinated assignment for the join task. But the unpredictability of actual runtimes and object sizes can create situations in which rigidly adhering to initial scheduling decisions could overload a worker. Accordingly, we adopted a two-step scheduling approach in which the initial ADFG can be adjusted during runtime. This turns out to be essential (see Section 6.3.1).
Navigator’s Task Dispatcher also keeps track of the tasks it has accumulated on its local execution queue, and estimates the wait time on this worker for a newly arriving task based on its current queuing situation. Task Dispatcher disseminates this information to all other workers through Global State Monitor. Schedulers on other servers could use this information to produce task assignments.
@@ 3.3 GPU Memory Manager
The purpose of this component is to manage ML models on worker nodes: the Navigator cache and the GPU model execution memory. Navigator cache holds reusable model objects in compressed form, whereas the execution memory is used to hold input objects for tasks and a decompressed model for each currently-active task.
As mentioned in Sections 1 and 2.2, GPU memories will often be too small to simply hold copies of every ML model that might be needed, and yet it is costly to fetch large models at the last instant when a task will be executed. Fortunately, ML models are reusable across different instances of the same pipeline and even across different pipelines, as we saw in Figure 1 where a single model was able to produce output in multiple languages. Retention of models in GPU memory can thus have a significant impact on job completion times. Prior work has often employed a static model placement, designating some set of workers that will "serve" each model, and managing GPU memory using a policy that the ML logic itself ultimately controls. In contrast, we take the approach of scheduler-triggered GPU memory management, in which the scheduler decides which worker will run each task, at which point the worker itself makes local decisions about model placement (both fetching and eviction) based on its assigned tasks. This design is especially beneficial for dynamic workloads with high variation or a sudden burst of requests for the same pipeline, because the Navigator scheduler can dynamically add additional workers to accommodate bursts of demand for particular MLs.
In the experiments in Section 6, we can see that proper task assignment depends upon careful management of model placement, and that Navigator is able to adjust to workloads with variations. While Navigator’s fetching and eviction policies are configurable, an important design element is to look ahead in the execution queue and consider what policies will be required for the currently queued tasks. We number active ML models in a small id space (currently 0..63), enabling the GPU Memory Manager to publish a bitmap encoding the current contents of the Navigator cache for each worker. The effect is that all workers know the cache contents of all their peers and can take this into consideration when making task assignments, thereby opportunistically reusing models already in GPU memory.
@@ 3.4 Global State Monitor
Navigator’s Global State Monitor manages a distributed table of system state. It runs on all workers and holds per-worker queue processing time and GPU memory (cache) contents. The queue processing time is the time to complete all tasks currently on the worker’s execution queue. The Navigator cache information includes the bitmap specifying which models are in the cache, as well as the amount of free memory in the cache.
To distribute updates, we take advantage of a modern multicast protocol optimized to leverage fast networking [37]. Each worker periodically multicasts updates about its local worker’s state information to all other workers. However, we limit the frequency of these updates to ensure that overhead will be low. The staleness of the information a worker sees about other workers has an upper bound that is equal to the dissemination interval. In the experiments at Section 6, we showed the scheduling mechanism has a high tolerance of information staleness, and enabled us to configure the update frequency to ensure that state information remains sufficiently accurate for Navigator’s task assignment decisions.
4 Navigator Scheduler Design
Navigator employs a heuristic scheduling mechanism with dynamic coordination between workers. The mechanism exhibits an important difference from schedulers in prior literature in that Navigator manages GPU memory and balances task assignment using a metric that reflects ML model object placements. The experiments reported in Section 6 confirm that this balance significantly improves end-to-end latency for job instances.
@@ 4.1 Parameters
The Navigator algorithm uses static information from the job DFG, the ADFG data produced by the initial scheduler, information from the repository of workflow profiles, and real-time information about worker states provided by Global State Monitor.
@@ 4.2 Planning Phase
The task assignment algorithm used during Navigator’s job planning phase maps tasks in a job instance to workers. This algorithm is inspired by a well-researched process scheduling algorithm, Heterogeneous Earliest Finish Time (HEFT) [8]. Like HEFT, Navigator uses an upward ranking to prioritize the task scheduling order and selects the workers with the earliest start time. However, Navigator extends the classic HEFT in several respects. HEFT does not consider the load of a machine when performing the scheduling, which in our setting could result in inefficient task assignment. Additionally, HEFT lacks mechanisms to factor in the potential benefits of co-location of the task with cached model objects on which it depends, or with input data. A third issue arises because HEFT locks down a plan for the entire job at the outset, preventing adjustments in the event that worker state changes in a significant way as the job instance proceeds. We address the first two limitations in the Navigator’s job planning algorithm and the third in its dynamic adjustment phase.
@@ 4.2.1 Vertex Ranking
Like HEFT, Navigator starts by assigning a rank to each task $t$. We use $t\prec t^{\prime}$ to denote that task $t$ is a direct predecessor of $t^{\prime}$ in the DFG. Since the target worker $w$ is unknown at the time of ranking, $R(t)$ is the average of $R(t,w)$ over the worker set. The rank is defined recursively as
@@
Notice that many aspects of task ranking can be performed statically. Navigator carries out these computations just once, when the DFG is initially loaded, then saves the results into its repository. Later, when dynamic inputs become available, the statically-computed rankings will merely need to be updated, not recomputed from scratch.
@@ 4.2.2 Task Assignments
The role of Navigator’s job instance planning phase is to produce an ADFG, which is a map from task ID to worker ID. The algorithm starts by considering the vertex rankings associated with the DFG. Navigator’s Scheduler module iterates through the tasks in descending order of priority, carrying out per-task scheduling decisions.
@@
$\emph{TD}_{model}(t,w)$ is the time for worker $w$ to load ML model $m$ for use by task $t$. Leveraging the data managed by the Global State Monitor, the scheduler can determine which GPUs currently hold decompressed instances of each of the models in the Navigator cache for all workers in the system. If a task is assigned to a worker that will need to fetch an ML model, the required time is computed as
@@
$\begin{cases}0,\textbf{ if}\text{ $m_{t}\in$ \emph{Navigator cache} on $w$}\\ \emph{TD}_{\emph{model}}(m,w),\textbf{ if}\text{ $m_{t}\not\in$ cache}\land|m_{t}|\leq\emph{AVC}(w)\\ \emph{TD}_{\emph{model}}(m,w)+\emph{eviction penalty},\textbf{ otherwise}\end{cases}$ (2)
Eviction penalty. Consider some task $t$ that needs to be scheduled. Navigator could schedule $t$ on a worker $w$ that already holds needed model $m$, or could also expand the pool of workers by assigning to some other worker $w^{\prime}$ where $m$ is not currently in cache. Naively, we should prioritize $w$ over $w^{\prime}$ if the expected delay at $w$ is longer than the transfer delay for loading $m$ at $w^{\prime}$ but this overlooks the likelihood that some other model $m^{\prime}$ will be evicted from cache by $w^{\prime}$ to make room for $m$. The penalty is intended to capture this effect.
@@ 5 Implementation
Navigator runs as a component of an open-source platform called Cascade [68, 37], which was created to reduce overheads for ML jobs. Prior to our work, Cascade held data (like a key-value store, file system, or message queuing middleware system). It also offered a hosting environment for low-latency ML tasks. Navigator enriches this with ML scheduling.
@@
Cascade additionally includes a storage/retrieval layer for in-memory replication of small data objects. This layer is called the shared state table (SST), and is used by Navigator to disseminate scheduler metadata (see Section 5.2).
Lacking Navigator, Cascade offered an automated hash-based load-balancer that was used in situations where a pool of servers all could handle a given request. Integration of Navigator with Cascade centered on replacing this mechanism with the decentralized Navigator scheduler and extending Cascade’s preexisting support for GPU accelerators with Navigator’s GPU memory management functionality.

### 28 Feb 2024 v2

# Compass: A Decentralized Scheduler for Latency-Sensitive ML Workflows

@@ Abstract
We consider ML query processing in distributed systems where GPU-enabled workers coordinate to execute complex queries: a computing style often seen in applications that interact with users in support of image processing and natural language processing. In such systems, coscheduling of GPU memory management and task placement represents a promising opportunity. We propose Compass, a novel framework that unifies these functions to reduce job latency while using resources efficiently, placing tasks where data dependencies will be satisfied, collocating tasks from the same job (when this will not overload the host or its GPU), and efficiently managing GPU memory. Comparison with other state of the art schedulers shows a significant reduction in completion times while requiring the same amount or even fewer resources. In one case, just half the servers were needed for processing the same workload.
@@ 1 Introduction
The work described here is motivated by a recent trend: under pressure to respond rapidly to incoming events, the cloud is expanding to include edge clusters physically close to the end-user. Our new scheduler, Compass, targets this setting, managing resources in ways that improve job completion delay with the same (or reduced) level of resources. Machine intelligence is already popular in interactive applications, and projected to gain importance for at least another decade, so progress could be broadly valuable.
@@
Our work represents edge ML applications as acyclic data-flow graphs (DFGs). Each ML instance runs in its own address space hence a single application could include ML pipelines implemented using different tools (we currently support PyTorch [34], TensorFlow [19], and MXNet [45]). A triggering event launches one job, and each DFG step within it runs as a task on a single server (we will relax these restrictions in future work). Compass plays two roles: platform-level GPU cache management and job/task placement. Our algorithm has the primary goal of minimizing job latency, with secondary goals of achieving efficient hardware utilization and packing work into as few servers as feasible so that unused hardware can be put into a power-saving mode.
An important decision was to make Compass fully decentralized and symmetric: every worker node can schedule tasks on every other worker. While many existing schedulers [36, 53, 50, 47, 64, 48] are centralized, permitting fully coordinated decision-making, centralization introduces overheads due to the need to consult the scheduler and to notify it of changing conditions, creating possible bottlenecks. Early in our effort, we realized that in edge systems task execution times are not fully predictable, hence worker loads and GPU memory contents can quickly evolve in unanticipated ways. Decentralized tracking of system state was traditionally assumed to be far more costly, but modern data replication is inexpensive and scalable [37]. Our decentralized solution has low overheads and outperforms centralized alternatives.
@@
    Experiments and simulation showing that Compass can reduce latency by 2x-6x. Compass never consumes more resources than existing schedulers and requires as few as half the servers to run an identical workload.
@@ 2.1 Dataflow Graphs
Compass’s DFG representation is similar to that used in prior work [10, 16, 74, 53]. As noted earlier, these are directed acyclic graphs $G=(E,V)$, where vertex $v\in V$ represents an ML computation. Edge $e\in E$ represents precedence constraints: output from the upstream ML will become input to the downstream ML. We attach a diamond box to vertex $v$ to represent data dependencies: the ML model and other objects that $v$ requires to perform its computational task. As seen in Figure 1, identical colors denote the same ML model, while different ones denote different models. Our figures omit details such as object sizes and estimated job/task execution times, but the DFG also includes them. Our experimental section used these workflows as its scheduling target.
@@ 2.2 Deployment Assumptions
Compass could be used in a cloud datacenter, but will often be deployed in edge clusters located near data-capture devices, an approach also adopted in prior work  [69]. These clusters would be more limited than full-fledged cloud servers but still could host multiple DFGs to serve mixes of edge requests. Using secure cloud links, edge applications would handle time-sensitive tasks while drawing on the full cloud for aspects off the low-latency critical path.
@@ 2.3 Scheduler Objectives
Our four example ML workflows all require fast responses. For example, when performing language translation and image generation, end-to-end performance of the DFG determines the wait time the human user will perceive. Compass’s primary role is to minimize end-to-end latency, defined as the difference between the end time (when the processing of the exit task is finished) and the start time (when the job instance is generated). Minimizing latency in DFG structured workflows is NP-complete, hence Compass will inevitably depend on heuristics [9, 11, 7]. We do not explicitly optimize for resource consumption or energy, but our algorithm turns out to be parsimonious in both respects.
@@ 3 System Architecture
As shown in Figure 2, we assume that each worker consists of a host supporting storage, host computation, and a GPU unit for heavy lifting. All workers run Compass, which is a fully decentralized and symmetric: every worker has visibility of the system state and can assign tasks to any of its peers.
Figure 2: Worker Components in Compass
@@
The architecture of Compass is seen in Figure 2 and consists of the following components: Workflow Profiling collects profiled information about the vertices in a DFG such as average execution times. The Decentralized Global State Monitor collects and periodically exchanges information about all the workers: their load, cache bitmap, etc. Compass’s Scheduler component creates a ADFG and later adjusts it. The task dispatcher executes the decisions taken by the Scheduler, sending task start requests to the correct worker as the prior task(s) complete. The GPU Memory Manager fetches models on the fly and manages a GPU cache. Task processing is managed by the Execution Engine, which includes a plug-in customized for each supported ML framework. We now describe each of the components, while the details of Compass’s scheduling algorithm are presented in Section 4.
@@ 3.1 Repository of Workflow Profiles
Compass’s Workflow Profiles Repository holds meta-information about DFGs. In addition to static information (expected runtime costs and input/output object sizes), each task is annotated with the actual sizes of known inputs.
@@ 3.2 Scheduler and Task Dispatcher
Preliminary scheduling decisions enable Compass to collocate a series of tasks on the same worker. This saves time by avoiding the need to move the output of one task to the location where the next worker would run. Moreover some aspects of task assignment can only occur at the outset: if a set of tasks are followed by a join, they would have no way to make a coordinated assignment for the join task. But the unpredictability of actual runtimes and object sizes can create situations in which rigidly adhering to initial scheduling decisions could overload a worker. Accordingly, we adopted a two-step scheduling approach in which the initial ADFG can be adjusted during runtime. This turns out to be essential (see Section 6.3.1).
Compass’s Task Dispatcher also keeps track of the tasks it has accumulated on its local execution queue, and estimates the wait time on this worker for a newly arriving task based on its current queuing situation. Task Dispatcher disseminates this information to all other workers through Global State Monitor. Schedulers on other servers could use this information to produce task assignments.
@@ 3.3 GPU Memory Manager
The purpose of this component is to manage ML models on worker nodes: the Compass cache and the GPU model execution memory. Compass cache holds reusable model objects in compressed form, whereas the execution memory is used to hold input objects for tasks and a decompressed model for each currently-active task.
As mentioned in Sections 1 and 2.2, GPU memories will often be too small to simply hold copies of every ML model that might be needed, and yet it is costly to fetch large models at the last instant when a task will be executed. Fortunately, ML models are reusable across different instances of the same pipeline and even across different pipelines, as we saw in Figure 1 where a single model was able to produce output in multiple languages. Retention of models in GPU memory can thus have a significant impact on job completion times. Prior work has often employed a static model placement, designating some set of workers that will "serve" each model, and managing GPU memory using a policy that the ML logic itself ultimately controls. In contrast, we take the approach of scheduler-triggered GPU memory management, in which the scheduler decides which worker will run each task, at which point the worker itself makes local decisions about model placement (both fetching and eviction) based on its assigned tasks. This design is especially beneficial for dynamic workloads with high variation or a sudden burst of requests for the same pipeline, because the Compass scheduler can dynamically add additional workers to accommodate bursts of demand for particular MLs.
In the experiments in Section 6, we can see that proper task assignment depends upon careful management of model placement, and that Compass is able to adjust to workloads with variations. While Compass’s fetching and eviction policies are configurable, an important design element is to look ahead in the execution queue and consider what policies will be required for the currently queued tasks. We number active ML models in a small id space (currently 0..63), enabling the GPU Memory Manager to publish a bitmap encoding the current contents of the Compass cache for each worker. The effect is that all workers know the cache contents of all their peers and can take this into consideration when making task assignments, thereby opportunistically reusing models already in GPU memory.
@@ 3.4 Global State Monitor
Compass’s Global State Monitor manages a distributed table of system state. It runs on all workers and holds per-worker queue processing time and GPU memory (cache) contents. The queue processing time is the time to complete all tasks currently on the worker’s execution queue. The Compass cache information includes the bitmap specifying which models are in the cache, as well as the amount of free memory in the cache.
To distribute updates, we take advantage of a modern multicast protocol optimized to leverage fast networking [37]. Each worker periodically multicasts updates about its local worker’s state information to all other workers. However, we limit the frequency of these updates to ensure that overhead will be low. The staleness of the information a worker sees about other workers has an upper bound that is equal to the dissemination interval. In the experiments at Section 6, we showed the scheduling mechanism has a high tolerance of information staleness, and enabled us to configure the update frequency to ensure that state information remains sufficiently accurate for Compass’s task assignment decisions.
4 Compass Scheduler Design
Compass employs a heuristic scheduling mechanism with dynamic coordination between workers. The mechanism exhibits an important difference from schedulers in prior literature in that Compass manages GPU memory and balances task assignment using a metric that reflects ML model object placements. The experiments reported in Section 6 confirm that this balance significantly improves end-to-end latency for job instances.
@@ 4.1 Parameters
The Compass algorithm uses static information from the job DFG, the ADFG data produced by the initial scheduler, information from the repository of workflow profiles, and real-time information about worker states provided by Global State Monitor.
@@ 4.2 Planning Phase
The task assignment algorithm used during Compass’s job planning phase maps tasks in a job instance to workers. This algorithm is inspired by a well-researched process scheduling algorithm, Heterogeneous Earliest Finish Time (HEFT) [8]. Like HEFT, Compass uses an upward ranking to prioritize the task scheduling order and selects the workers with the earliest start time. However, Compass extends the classic HEFT in several respects. HEFT does not consider the load of a machine when performing the scheduling, which in our setting could result in inefficient task assignment. Additionally, HEFT lacks mechanisms to factor in the potential benefits of co-location of the task with cached model objects on which it depends, or with input data. A third issue arises because HEFT locks down a plan for the entire job at the outset, preventing adjustments in the event that worker state changes in a significant way as the job instance proceeds. We address the first two limitations in the Compass’s job planning algorithm and the third in its dynamic adjustment phase.
@@ 4.2.1 Vertex Ranking
Like HEFT, Compass starts by assigning a rank to each task $t$. We use $t\prec t^{\prime}$ to denote that task $t$ is a direct predecessor of $t^{\prime}$ in the DFG. Since the target worker $w$ is unknown at the time of ranking, $R(t)$ is the average of $R(t,w)$ over the worker set. The rank is defined recursively as
@@
Notice that many aspects of task ranking can be performed statically. Compass carries out these computations just once, when the DFG is initially loaded, then saves the results into its repository. Later, when dynamic inputs become available, the statically-computed rankings will merely need to be updated, not recomputed from scratch.
@@ 4.2.2 Task Assignments
The role of Compass’s job instance planning phase is to produce an ADFG, which is a map from task ID to worker ID. The algorithm starts by considering the vertex rankings associated with the DFG. Compass’s Scheduler module iterates through the tasks in descending order of priority, carrying out per-task scheduling decisions.
@@
$\emph{TD}_{model}(t,w)$ is the time for worker $w$ to load ML model $m$ for use by task $t$. Leveraging the data managed by the Global State Monitor, the scheduler can determine which GPUs currently hold decompressed instances of each of the models in the Compass cache for all workers in the system. If a task is assigned to a worker that will need to fetch an ML model, the required time is computed as
@@
$\begin{cases}0,\textbf{ if}\text{ $m_{t}\in$ \emph{Compass cache}on $w$}\\ \emph{TD}_{\emph{model}}(m,w),\textbf{ if}\text{ $m_{t}\not\in$ cache}\land|m_{t}|\leq\emph{AVC}(w)\\ \emph{TD}_{\emph{model}}(m,w)+\emph{eviction penalty},\textbf{ otherwise}\end{cases}$ (2)
Eviction penalty. Consider some task $t$ that needs to be scheduled. Compass could schedule $t$ on a worker $w$ that already holds needed model $m$, or could also expand the pool of workers by assigning to some other worker $w^{\prime}$ where $m$ is not currently in cache. Naively, we should prioritize $w$ over $w^{\prime}$ if the expected delay at $w$ is longer than the transfer delay for loading $m$ at $w^{\prime}$ but this overlooks the likelihood that some other model $m^{\prime}$ will be evicted from cache by $w^{\prime}$ to make room for $m$. The penalty is intended to capture this effect.
@@ 5 Implementation
Compass runs as a component of an open-source platform called Cascade [68, 37], which was created to reduce overheads for ML jobs. Prior to our work, Cascade held data (like a key-value store, file system, or message queuing middleware system). It also offered a hosting environment for low-latency ML tasks. Compass enriches this with ML scheduling.
@@
Cascade additionally includes a storage/retrieval layer for in-memory replication of small data objects. This layer is called the shared state table (SST), and is used by Compass to disseminate scheduler metadata (see Section 5.2).
Lacking Compass, Cascade offered an automated hash-based load-balancer that was used in situations where a pool of servers all could handle a given request. Integration of Compass with Cascade centered on replacing this mechanism with the decentralized Compass scheduler and extending Cascade’s preexisting support for GPU accelerators with Compass’s GPU memory management functionality.